Two terms of an arithmetic sequence are a9 = 85 and a30 = 295. Find a rule for the nth term.
Answer by scott8148(6628)

the 21 differences between a9 and a30 equals 210 (295-85)
so the difference between consecutive terms is 10 (210/21)

an = nd + a0 ___ a9 = 9(10) + a0 ___ 85 = 90 + a0 ___ -5 = a0

an = n(10) - 5